OpenStudy (anonymous):

solve the equation using square roots. x^2+20=4

terenzreignz (terenzreignz):

First, you'll want to get all the constants to the right-side. Can you bring the +20 to the right?

OpenStudy (anonymous):

im so confused can u just help me with the answer

OpenStudy (tkhunny):

Calm. There si something important to learn, here. The "answer" is what you learn. There isn't a number to write or a box to check. Subtract 20 from both sides and see what happens.

OpenStudy (tanya123):

\(x^2+20 = 4\) \(x^2+16 = 0\)

terenzreignz (terenzreignz):

@Compound Confusion is a truth of Maths study (unfortunately) but I'm going to need you to try all the same :) As previously stated, you can get rid of the +20 on the left-side by subtracting 20 from both sides. Leaving you with...

OpenStudy (tkhunny):

I like that. Once you see \(x^{2} + 16\), alarms should go off in your head!! "Hey, I can't factor that! It's a SUM of squares!" However, we are to solve using square roots.

OpenStudy (anonymous):

x^2 +16=0 x^2 = -16 x = +/- 4i, where i is a complex number equal to sqrt (-1)